OpenStudy (hkmiu):

x^2+y^2=2y in polar form

OpenStudy (dumbcow):

\[x^2 + y^2 = r^2\] \[y = r \sin \theta\] therefore \[\rightarrow r^2 = 2r \sin \theta\] \[\rightarrow r = 2 \sin \theta\]

OpenStudy (hkmiu):

thanks for the explanation